The height of the Eiffel Tower is 984 ft. A souvenir model of the tower is 6 in. tall. At 5 P.M. in Paris, the shadow of the souvenir model is 8 in. long. The Eiffel Tower and its shadow determine two legs of a right triangle that are similar to the two legs of a right triangle determined by the souvenir model and its shadow. About how long is the shadow of the Eiffel Tower?
Answer by stanbon(48568)

The height of the Eiffel Tower is 984 ft. A souvenir model of the tower is 6 in. tall. At 5 P.M. in Paris, the shadow of the souvenir model is 8 in. long. The Eiffel Tower and its shadow determine two legs of a right triangle that are
similar to the two legs of a right triangle determined by the souvenir model and its shadow. About how long is the shadow of the Eiffel Tower?
-------------------
6 in / 8 in = 984 ft/ x ft
-------------------------
3/4 = 984/x
x/4 = 984/3
x = 4*(984/3)
x = 1312 ft. (length of shadow of Eiffel Tower at 5 P.M.)
==================
Cheers,
Stan H.